# Trustworthy Answers, Messier Data: Bridging the Gap in Low-Resource Retrieval-Augmented Generation for Domain Expert Systems

### v2

@@ 4.3.1 Multi-modal Specific Questions
20 | 0.4949 | 0.5923 | 0.5995 | |||
30 | 0.4933 | 0.5965 | 0.6041 |
4.4 Analysis
Figure 3 illustrates the trade-off between retrieval success and re-ranking performance as the number of retrieved chunks increases. While increasing improves the retrieval success rate (e.g., from 92% at = 10 to 96% at = 30), it can lead to diminishing improvements in re-ranking performance (e.g., MAP@5 decreases from 0.74 at = 10 to 0.73 at = 30). This result highlights the importance of carefully selecting to balance retrieval success and re-ranking quality, as overly increasing may not yield proportional benefits for re-ranking.
An analysis of the failed cases during the retrieval and ranking phase identified two error types: (a) Top-10 retrieval failure (23.3%) and (b) Retrieval success but incorrect re-ranker top-1 (76.7%). Type (a) were mainly caused by the open-ended nature of the questions, which led to multiple relevant documents beyond the gold context (Figure 18 in A.3.2). In contrast, type (b) occurred with more specific and detailed questions, where there were several relevant chunks from the same vehicle crash collision test, in addition to the gold chunk, making it difficult to identify a single correct one (Figure 19 in A.3.2). Failures in both types were often due to the presence of multiple valid chunks, rather than the ranking of irrelevant chunks. Table 4 presents a human evaluation of 100 sampled cases for each error type, assessing whether the re-ranker top-1 was relevant to the given question, even though it did not match the gold chunk. A screenshot of the human evaluation interface is in Figure 21 (A.5).
5 Experiment 2: Answer Generation
5.1 Models
We use Qwen-2.5 Yang et al. (2024) as our backbone LLMs, one of the few multilingual models that officially supports Korean. Notably, there are currently no Korean-centric open-source LLMs with sufficient context length to address our requirements. We fine-tune Qwen 2.5 (14B, 72B) on our training dataset, performing full fine-tuning on the 14B model and applying Low-Rank Adaptation (LoRA) (Hu et al., 2021) for the 72B model. Experimental details including hyperparameters and GPU configurations, are in A.1.2.
5.2 Evaluation
5.2.1 LLM-as-a-Judge
As illustrated in Figure 5, we employ the LLM-as-a-judge approach (Zheng et al., 2023) to evaluate the performance of both LLM-only and RAG models. For the LLM evaluation, we compare four model variations: the vanilla and fine-tuned versions of 14B and 72B models. GPT-4o444https://platform.openai.com/docs/models/gpt-4o is used to rank anonymized models (A, B, C, and D) based on factual correctness, helpfulness, and informativeness. These metrics are chosen to evaluate distinct, non-overlapping dimensions:
@@ 5.2.1 LLM-as-a-Judge
For RAG evaluation, we use the best-performing one among four models and compare its performance with and without RAG integration. Instead of pairwise comparisons, we employ single-answer evaluations, scoring responses (1–5) on three aspects. This approach offers a more detailed assessment, emphasizing how RAG impacts specific aspects of response quality. All the evaluation prompts used are detailed in A.2.2.
5.2.2 Human Evaluation
To assess the reliability of the RAG evaluation using the LLM-as-a-judge method, we conducted a human evaluation to compare its alignment with the LLM evaluation. An expert evaluator, a native Korean automotive engineer, assessed 100 randomly selected and anonymized responses. These responses were generated by the 72B fine-tuned model, both with and without RAG integration. A screenshot of the human evaluation interface used is in shown Figure 22 (A.5).
5.3 Result
5.3.1 LLM-only Comparison
@@
5.3.3 Comparison with Human Evaluation
Table 6 shows that human evaluation aligns with GPT evaluation, both assigning higher scores to the RAG model across all three aspects. Figure 8 visualizes the differences in scoring tendencies between human and GPT evaluations. Notably, correctness and informativeness show an inverted pattern: human favors more polarized correctness scores (1 and 5), while GPT prefers mid-range (2 and 3). Additionally, human assigns higher helpfulness scores more frequently, as seen in the positive values for 4 and 5. While differences in score distributions exist between evaluators, these differences are independent of the specific models. Both evaluations consistently show that RAG outperforms non-RAG by achieving higher scores.
@@ 6 Experiment 3: Reference Matching
Segment 5 | Sentence 8 | All of these improvements aim to optimize the body structure and prevent component fracture during collisions. | Chunk 1,2,3 |
6.1 Evaluation
To evaluate the reference matching task, we randomly sample 100 triplets () from the test set and perform human annotation to identify which chunk () each sentence in the generated answer () referenced. Each sentence is annotated with one or more chunks, allowing for multiple references. The interface of the annotation tool is in Figure 23 (A.5). We use sentence-level precision as the evaluation metric, defined as the proportion of sentences where the model prediction is included in the annotated reference set for an answer, averaged over the entire evaluation set. Since references are identified for concatenated sentences, selecting the best match for the grouped information is key. Recall would unfairly penalize valid predictions for not selecting every reference.
6.2 Result
Table 7 compares the reference matching performance of our algorithm (Section 3.4) without thresholding and the fine-tuned Qwen 72B, while Figure 10 illustrates the score distribution and the increase in precision with score thresholding for our algorithm. Without thresholding, our algorithm achieves a sentence-level precision of 0.72, while the LLM reaches 0.81, showing a noticeable gap in performance. However, when thresholding is applied, the precision reaches 0.86 at a threshold of 0.5, with performance improving proportionally as the threshold is increased further. This demonstrates the reliability of our algorithm’s scoring, allowing for controlled adjustment of reference matching quality through threshold selection. Our re-ranker used in the matching algorithm has only 0.5B parameters, compared to the LLM’s 72B, demonstrating that our algorithm can achieve impressive performance with a smaller model and faster inference speed. The LLM prompt used for reference matching is in Figure 17 (A.2.3). The answers are divided into 2-3 segments on average, and we identify two types of reference matching: one where all the necessary information to answer the question is contained within a single chunk, resulting in a single-segment answer, and another where information from multiple chunks is needed to answer the question, leading to a multi-segment answer. The first type often occurs in factual questions related to specific vehicle crash collision test, while the second type is more common in open-ended questions that require referencing multiple tests (Figure 9).
@@ 6.2 Result
3.92 | ||||||
LLM |
13.54 |
7 Conclusion
In this work, we presented a low-resource approach to RAG for domain expert QA, utilizing heterogeneous data sources. We showed that our data generation pipeline, guided by expert-driven prompting, facilitates an effective RAG system by converting multi-modal raw data. By enhancing the retrieval phase with re-ranker, we demonstrated that the final answers are both accurate and traceable to relevant sources. These findings highlight the potential of RAG in data-scarce domains and its applicability to specialized knowledge-intensive tasks. As future work, we are exploring open-source Vision models to better handle multi-modal questions, addressing current limitations discussed in Section 8.
8 Limitations
Constraints in Backbone Model Selection The selection of a backbone model was constrained by the need for an open-source model supporting Korean with sufficient context length, limiting the available options across retrieval and LLM components. BGE-M3 for retriever & re-ranker, along with Qwen2.5 for the LLM, were the options available to us. However, it is worth noting that the newly released Qwen2.5-1M Yang et al. (2025) offers a longer context length, and our preliminary experiments demonstrated its potential for better handling long inputs. As the availability of open-source LLMs continues to grow, we anticipate that this limitation will gradually be resolved over time.
Restricted Use of API-based LLM for Initial Data Processing While our ultimate goal is to develop a fully local system, we initially relied on the API-based LLM, specifically the Claude API, for limited data processing. Among the available open-source LLMs, none met the requirements due to context limits and certain performance considerations, making them unsuitable for our needs. We determined that leveraging an API-based LLM with expert-driven prompting would be more effective in terms of both quality and cost compared to full human annotation. However, this reliance was limited to the initial data required for fine-tuning in the open-source model, and any future updates to the document pool will completely eliminate the need for external dependencies.
Weakness in Multi-modal Questions Our RAG system exhibited relatively lower performance when retrieving contexts for multi-modal specific questions that require information from tables, graphs, and images to answer the given question (Section 4.3.1). We hypothesize that this is due to the loss of rich content when multi-modal elements are converted into plain text during the data processing stage. To address this, we are experimenting with Llama 3.2 Vision555https://huggingface.co/meta-llama/Llama-3.2-11B-Vision to improve the conversion of multi-modal elements into text, aiming for better representation of these elements.
Experiments Limited by Computational Constraints The number of experiments we could conduct on the generation model was constrained by computational limitations. Fine-tuning the model on instead of (Section 3.3) proved impractical due to the extensive context length, which exponentially increased GPU memory requirements, surpassing our available computational resources. Although we observed that Qwen models inherently possess summarization capabilities for the given , and fine-tuning on yielded effective results, we were unable to explore a variety of learning strategies for the generation model within the RAG framework.
Appendix A Appendix
A.1 Experimental Details
A.1.1 Dataset
We split the generated dataset at the chunk level, slide level for slides, and chapter level for the textbook, using an 8:1:1 ratio for training, validation, and testing. Table 8 summarizes the dataset splits, while token length statistics for the training data are presented in Table 9.
Source | Data | Train | Val | Test | Total |
Test Report | Chunk | 3,729 | 466 | 467 | 4,662 |
Q&A Pair | 47,660 | 5,823 | 5,919 | 59,402 | |
Meeting Report | Chunk | 705 | 88 | 89 | 882 |
Q&A Pair | 6,144 | 752 | 800 | 7,696 | |
Textbook | Chunk | 64 | 8 | 9 | 81 |
Q&A Pair | 1,182 | 162 | 161 | 1,505 |
Min | Mean | Max | |
Question | 9 | 35 | 82 |
Answer | 6 | 56 | 271 |
Chunk (Ver.0) | 66 | 843 | 9,528 |
Chunk (Ver.1) | 106 | 884 | 9,528 |
A.1.2 Model Training
The retriever was trained for 10 epochs, determined to be optimal, using 4 48GB RTX A6000 GPUs. The hyperparameters were set to the default configuration provided in the open-source implementation of BGE-M3. For the ranker, early stopping was employed by validating the model every 1,000 steps. Training was stopped if performance did not improve after 3 consecutive validations, and the best-performing model was selected. This training used 1 48GB RTX A6000 GPU with a learning rate of 1e-5, the AdamW optimizer Loshchilov and Hutter (2017), and a batch size of 4. For both components, validation was based on the MAP@10 metric evaluated on the validation set. To train our answer generation model, full fine-tuning was conducted for the 14B model with a batch size of 2, gradient accumulation steps of 64, a learning rate of 2e-5, and 3 training epochs using 3 80GB H100 GPUs. For the 72B model, LoRA was applied with a batch size of 1, gradient accumulation steps of 64, a learning rate of 2e-5, and 3 training epochs using 2 × 141GB H200 GPUs.
A.2 Prompt Engineering
A.2.1 Prompts for Data Generation
The prompt used to convert slides into Markdown text is shown in Figure 12, and the prompt for Q&A generation is depicted in Figure 13. An example of a generated Q&A pair from a chunk is shown in Figure 14. An average of twelve Q&A pairs were generated from a single chunk.
A.2.2 Prompts for Final Answer Evaluation
Our prompts for evaluating final answers from LLM-only models and comparing LLM-only and RAG models are in Figure 15 and 16, respectively.
A.2.3 Prompts for Reference Matching
The prompt used for LLM reference matching is shown in Figure 17.
A.3 Further Analysis
A.3.1 TS+TCL Effect on Ranker Training
Table 10 shows the re-ranking performance without TS+TCL training. When compared to the performance with TS+TCL training (Table 3), there is a noticeable decline across all metrics.
MAP@1 | MAP@5 | MAP@10 | Success@5 | |||||||||
10 |
20 |
30 |
A.3.2 Analysis of Retrieval & Ranking Phase
Figure 18 and 19 present real-world examples of two error types: (a) cases where the top-retrieval process failed, and (b) cases where retrieval succeeded but the ranker’s top-1 result did not match the gold context.
Question: Were there cases where the body rating significantly decreased? | |||||||||||||||||||||||||||||||
Model Top-1 Chunk | Gold Chunk | ||||||||||||||||||||||||||||||
Model Top-1 Chunk | Gold Chunk | ||||||||||||||||||||||||||
A.3.3 LLM-only vs. RAG
Table 11 presents detailed counts and score distributions from GPT-4o, comparing RAG and non-RAG.
Metric | Score | Non-RAG | RAG |
Correctness | 1 | 1,640 | 388 |
2 | 3,166 | 579 | |
3 | 1,457 | 705 | |
4 | 430 | 1,359 | |
5 | 187 | 3,849 | |
Avg. Score | 2.1799 | 4.1195 | |
Helpfulness | 1 | 758 | 154 |
2 | 2,819 | 522 | |
3 | 2,422 | 868 | |
4 | 722 | 1,649 | |
5 | 159 | 3,687 | |
Avg. Score | 2.5211 | 4.1908 | |
Informativeness | 1 | 461 | 142 |
2 | 2,764 | 564 | |
3 | 2,789 | 1,609 | |
4 | 756 | 2,974 | |
5 | 110 | 1,591 | |
Avg. Score | 2.6061 | 3.7715 |
A.4 Analysis of Reference Matching
An example where only a single chunk needs to be referred to answer the question is in Figure 20.
Question: What was the material of the back beam used in the HDC P5 40kph slope 30 (RH) test? | ||||
Answer | Segment 1 | Sentence 1 | The material of the back beam used in the HDC P5 40kph slope 30 (RH) test was aluminum. | Chunk 1 |
Sentence 2 | This information can be found in the test conditions section. |
A.5 Human Evaluation Interface
The human evaluation interfaces for assessing the retrieval and re-ranking phase, as well as the reliability of the RAG evaluation with LLM, are shown in Figures 21 and 22, respectively. The human evaluation interface used for annotating the evaluation dataset to aseess the reference matching task is shown in Figure 23.

### v5

@@ 4.1.3. Multi-modal Specific Questions
| 20 | 0.4949 | 0.5923 | 0.5995 | |||
| 30 | 0.4933 | 0.5965 | 0.6041 |
4.1.4. Analysis
Figure 4 illustrates the trade-off between retrieval success and re-ranking performance as the number of retrieved chunks increases. While increasing improves the retrieval success rate (e.g., from 92% at = 10 to 96% at = 30), it can lead to diminishing improvements in re-ranking performance (e.g., MAP@5 decreases from 0.74 at = 10 to 0.73 at = 30). This result highlights the importance of carefully selecting to balance retrieval success and re-ranking quality, as overly increasing may not yield proportional benefits for re-ranking. In addition, an analysis of the failed cases during the retrieval and ranking phase identified two error types: (a) Top-10 retrieval failure (23.3%) and (b) Retrieval success but incorrect re-ranker top-1 (76.7%). Type (a) were mainly caused by the open-ended nature of the questions, which led to multiple relevant documents beyond the gold context. In contrast, type (b) occurred with more specific and detailed questions, where there were several relevant chunks from the same vehicle crash collision test, in addition to the gold chunk, making it difficult to identify a single correct one. Failures in both types were often due to the presence of multiple valid chunks, rather than to the ranking of irrelevant chunks. Table 5 presents a human evaluation of 100 sampled cases for each error type, assessing whether the re-ranker top-1 was relevant to the given question, even though it did not match the gold chunk.
4.2. Answer Generation
4.2.1. Evaluation
We employ the LLM-as-a-judge approach (Zheng et al., 2023) to evaluate the performance of both LLM-only and RAG models. For the LLM evaluation, we compare four model variations: the vanilla and fine-tuned versions of 14B and 72B models. GPT-4o 555https://platform.openai.com/docs/models/gpt-4o is used to rank anonymized models (A, B, C, and D) based on factual correctness, helpfulness, and informativeness. These metrics are chosen to evaluate distinct, non-overlapping dimensions:
@@ 4.2.1. Evaluation
For RAG evaluation, we use the best-performing one among four models and compare its performance with and without RAG integration. Instead of pairwise comparisons, we employ single-answer evaluations, scoring responses (1–5) on three aspects. This approach offers a more detailed assessment, emphasizing how RAG impacts specific aspects of response quality. To assess the reliability of LLM evaluation, we conducted a human evaluation by a domain expert in automotive safety. inter-annotator agreement (IAA) could not be calculated due to resource constraints, and the evaluation was limited to 100 randomly selected and anonymized samples; scaling to more annotators and larger sets will be part of future work.
4.2.2. Result
@@
Table 8 shows that human evaluation aligns with GPT evaluation, both assigning higher scores to the RAG model across all aspects. Figure 8 visualizes the differences in scoring tendencies between human and GPT evaluations. Notably, correctness and informativeness show an inverted pattern: human favors more polarized correctness scores (1 and 5), while GPT prefers mid-range (2 and 3). Additionally, human assigns higher helpfulness scores more frequently, as seen in the positive values for 4 and 5. While differences in score distributions exist between evaluators, these differences are independent of the specific models. Both evaluations consistently show that RAG outperforms non-RAG by achieving higher scores.
@@ 4.3. Reference Matching
| Segment 5 | Sentence 8 | All of these improvements aim to optimize the body structure and prevent component fracture during collisions. | Chunk 1,2,3 | ||
4.3.1. Evaluation
To evaluate the reference matching task, we randomly sample 100 triplets () from the test set and manually annotate to identify which chunk () each sentence in the generated answer () referenced. Each sentence is annotated with one or more chunks, allowing for multiple references. We use sentence-level precision as the evaluation metric, defined as the proportion of sentences where the model prediction is included in the annotated reference set for an answer, averaged over the entire evaluation set. Since references are identified at the level of concatenated sentences, selecting the best match for the grouped information becomes most important. Evaluating recall would unfairly penalize valid predictions that do not select every reference.
4.3.2. Result
Table 9 compares our reference matching (Section 3.3) performance without thresholding and the fine-tuned Qwen 72B 666Prompt used for Qwen 72B reference matching is like: You are an expert tasked with analyzing user questions, an answer split into sentences, and a list of related chunks. Your job is to identify the single most relevant chunk for each sentence based on sentence indices provided. Map the segments to relevant chunks., while Figure 10 illustrates the score distribution and the increase in precision with score thresholding for our algorithm. Without thresholding, our algorithm achieves a sentence-level precision of 0.72, while the LLM reaches 0.81, showing a noticeable gap in performance. However, when thresholding is applied, the precision reaches 0.86 at a threshold of 0.5, with performance improving proportionally as the threshold is increased further. This demonstrates the reliability of our algorithm, allowing for controlled adjustment of reference matching quality through threshold selection. Our re-ranker used in the matching algorithm has only 0.5B parameters, compared to the LLM’s 72B, demonstrating that our algorithm can achieve impressive performance with a smaller model and faster inference speed. The answers are divided into 2-3 segments on average, and we identify two types of reference matching: one where all the necessary information to answer the question is contained within a single chunk, resulting in a single-segment answer, and another where information from multiple chunks is needed to answer the question, leading to a multi-segment answer. The first type often occurs in factual questions related to specific vehicle crash collision test, while the second type is more common in open-ended questions that require referencing multiple tests (Figure 9).
@@ 4.3.2. Result
| Ours | 2.7 ( 1.5) | 0.72 ( 0.27) | 3.92 | ||||
| LLM | 1.9 ( 0.9) | 0.81 ( 0.26) | 13.54 |
5. Conclusion
In this study, we presented an on-premise RAG-QA system tailored for internal corporate documents, with a focus on the automotive domain. The system comprises a domain-adapted retriever, re-ranker, LLM, and reference matcher, and has proven effective through comprehensive evaluations and a pilot test with internal employees. Ongoing work addresses current limitations, including the limited availability of open-source LLMs that support Korean and long-context processing. We are experimenting with newly released models such as Qwen2.5-1M (Yang et al., 2025), which may serve as viable alternatives. To better handle multi-modal questions involving tables and figures, we are also exploring LLaMA 3.2 Vision (AI, 2024) to enhance the textual representation of visual content. We believe our RAG development pipeline offers a practical solution for organizations seeking to effectively utilize fragmented internal documents in decision-making processes. We released the code—including data processing and RAG components training—as well as all the prompts used: https://github.com/emorynlp/MultimodalRAG
Acknowledgements.
We gratefully acknowledge the support of Hyundai Motor Company. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of Hyundai Motor Company.GenAI Usage Disclosure
We used ChatGPT 777https://chatgpt.com/ solely for minor English grammar and style corrections after the authors had drafted the full text.